# Enhanced Feature Learning via Regularisation: Integrating Neural Networks and Kernel Methods

### v1

@@ Abstract
We propose a new method for feature learning and function estimation in supervised learning via regularised empirical risk minimisation. Our approach considers functions as expectations of Sobolev functions over all possible one-dimensional projections of the data. This framework is similar to kernel ridge regression, where the kernel is $\mathbb{E}_{w}(k^{(B)}(w^{\top}x,w^{\top}x^{\prime}))$, with $k^{(B)}(a,b):=\min(|a|,|b|)\mathds{1}_{ab>0}$ the Brownian kernel, and the distribution of the projections $w$ is learnt. This can also be viewed as an infinite-width one-hidden layer neural network, optimising the first layer’s weights through gradient descent and explicitly adjusting the non-linearity and weights of the second layer. We introduce an efficient computation method for the estimator, called Brownian Kernel Neural Network (BKerNN), using particles to approximate the expectation. The optimisation is principled due to the positive homogeneity of the Brownian kernel. Using Rademacher complexity, we show that BKerNN’s expected risk converges to the minimal risk with explicit high-probability rates of $O(\min((d/n)^{1/2},n^{-1/6}))$ (up to logarithmic factors). Numerical experiments confirm our optimisation intuitions, and BKerNN outperforms kernel ridge regression, and favourably compares to a one-hidden layer neural network with ReLU activations in various settings and real data sets.
@@ 1 Introduction
Multiple index models have been extensively studied, leading to various methods for estimating the feature space. Brillinger (2012) introduced the method of moments for Gaussian data and one feature, by using specific moments to eliminate the unknown function. For features of any dimension, several methods have been proposed. Sliced inverse regression (SIR) (Li, 1991) uses second-order moments to identify effective dimensions by slicing the response variable and finding linear combinations of predictors, while improvements have been proposed (Yang et al., 2017), these methods heavily rely on assumptions about the covariate distribution shape and prior knowledge of the distribution. Iterative improvements have been an interesting line of work (Dalalyan et al., 2008), while optimisation-based methods like local averaging minimise an objective function to estimate the subspace (Fukumizu et al., 2009; Xia et al., 2002). Despite their practical performance, particularly the MAVE method (Xia et al., 2002), the theoretical guarantees show exponential dependence on the original data dimension, making them less suitable for high-dimensional settings.
@@
Regularising by adding a penalty term to the empirical risk minimisation objective guides specific estimator behaviours. In the context of feature learning, Rosasco et al. (2013) used derivatives for regularisation in nonparametric models, focusing on variable selection. While their method reduces to classical regularisation techniques for linear functions, it faces limitations: functions depending on a single variable do not belong in the chosen RKHS, using derivatives at data points limits the exploitation of regularity, and there is no benefit from hidden variables in the misspecified case. An improvement over this method was studied for both feature learning and variable selection by Follain and Bach (2024), where a trace norm penalty (Koltchinskii et al., 2011; Giraud, 2014) on the derivatives was used for the feature learning case. However, the dependency on the dimension of the rate did not allow high-dimensional learning. We can justify the use of trace norm penalties by considering the structure of neural networks. Under the multiple-index model, the weights $w_{1},\ldots,w_{m}$ of the first layer are expected to lie in a low-rank subspace of rank at most $k$. However, directly enforcing a rank constraint is not practical for optimisation. Therefore, we could use a relaxation such as $\Omega(f)=\operatorname*{tr}\big{(}\big{(}\sum_{j=1}^{m}w_{j}w_{j}^{\top}\big{)}^{1/2}\big{)}$, which is the trace norm of a matrix containing the weights, to approximate the rank constraint effectively. However, there is still the issue of multiple constant terms for a single weight. We will see specialised penalties for feature learning for a different family of functions.
@@ 2.1 Custom Space of Functions
Definition 1 (Infinite-Width Function Space))
@@
:=$\mathcal{F}_{\infty}:=\left\{f\mid f(\cdot)=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot)\,{\rm d}\mu(w)\right\},$
@@
The function space $\mathcal{F}_{\infty}$ is inspired by infinite-width single hidden layer neural networks: with the addition of the intercept $c$, each function in this space can be seen as the integral of a linear part $w$ and a non-linearity $g_{w}$ over some probability distribution, as in Equation (1) where the non-linearity is $\eta\sigma(\cdot)$. Thus here the activation functions are learnt.
@@
We now consider regularised empirical risk minimisation starting with a basic penalty. Let
$\Omega_{0}(f)=\inf_{c\in\mathbb{R},(g_{w})_{w}\in\mathcal{H}^{\mathcal{S}^{d-1}},\mu\in\mathcal{P}(\mathcal{S}^{d-1})}\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}\,{\rm d}\mu(w),$ (2)
such that $f=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot)\,{\rm d}\mu(w)$, where $\|g_{w}\|_{\mathcal{H}}^{2}:=\int_{-\infty}^{+\infty}g_{w}^{\prime}(t)^{2}\,{\rm d}t$. This penalty enforces the regularity of the function and, because we use penalizations with non-squared norms, limits the number of non-zero $g_{w}$. While this penalty is not specifically aimed at feature learning, by limiting the number of non-zero particles, it indirectly promotes feature learning to some extent. This serves as a starting point, and we introduce more targeted penalties in Section 2.5 with a stronger feature learning behavior. For $f\in\mathcal{F}_{m}$ written as in Definition 2, the penalty simplifies to $\Omega_{0}(f)=\frac{1}{m}\sum_{j=1}^{m}\|g_{j}\|_{\mathcal{H}}$. The learning objective is thus defined as
@@
$\min_{c\in\mathbb{R},(g_{w})_{w}\in\mathcal{H}^{\mathcal{S}^{d-1}},\mu\in\mathcal{P}(\mathcal{S}^{d-1})}\frac{1}{n}\sum_{i=1}^{n}\ell\bigg{(}y_{i},c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}x_{i}){\rm d}\mu(w)\bigg{)}+\lambda\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w),$ (4)
@@
$\min_{c\in\mathbb{R},w_{1},\ldots,w_{m}\in\mathcal{S}^{d-1},g_{1},\ldots,g_{m}\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\ell\bigg{(}y_{i},c+\frac{1}{m}\sum_{j=1}^{m}g_{j}(w_{j}^{\top}x_{i})\bigg{)}+\lambda\frac{1}{m}\sum_{j=1}^{m}\|g_{j}\|_{\mathcal{H}}.$ (5)
@@ 2.2 Properties of Reproducing Kernel Hilbert Space $\mathcal{H}$ and Kernel $k$
with the inner product $\langle\tilde{g},g\rangle=\int_{\mathbb{R}}\tilde{g}^{\prime}g^{\prime}$. This space is a reproducing kernel Hilbert space with the reproducing kernel $k^{(B)}(a,b)=(|a|+|b|-|a-b|)/2=\min(|a|,|b|)\mathds{1}_{ab>0}$. This kernel, which can be referred to as the “Brownian” kernel, corresponds to the covariance of the Brownian motion at times $a$ and $b$ (Mishura and Shevchenko, 2017). Consequently, we have the reproducing property
$\forall g\in\mathcal{H},\forall a\in\mathbb{R},g(a)=\langle g,k^{(B)}_{a}\rangle,$
where $k^{(B)}_{a}:b\in\mathbb{R}\to k^{(B)}(a,b)\in\mathbb{R}$. As a reproducing kernel, it is positive definite, meaning that for any $n\in\mathbb{N}$, $\alpha\in\mathbb{R}^{n}$, and $a\in\mathbb{R}^{n}$, we have $\sum_{i,j=1}^{n}\alpha_{i}k^{(B)}(a_{i},a_{j})\alpha_{j}\geq 0$. Additionally, we observe that $\|k^{(B)}_{a}\|^{2}_{\mathcal{H}}=|a|$ and $\|k^{(B)}_{a}-k^{(B)}_{b}\|^{2}_{\mathcal{H}}=|a-b|$. It is also noteworthy that by definition, the functions in $\mathcal{H}$ are necessarily continuous, in fact even $1/2$-Hölder continuous as we see in Lemma 3.
@@ 2.3 Characterisation of $\mathcal{F}_{\infty}$
$\mathcal{F}_{\infty}$ is a vector space, $\max(f(0),\Omega_{0}(f))$ is a norm on $\{f\in\mathcal{F}_{\infty}\mid\Omega_{0}(f)<\infty\}$. For $f\in\mathcal{F}_{\infty}$ with $\Omega_{0}(f)<\infty$, the function $f$ is $1/2$-Hölder continuous with constant $\Omega_{0}(f)$, i.e., $|f(x)-f(x^{\prime})|\leq\Omega_{0}(f)\sqrt{\|x-x^{\prime}\|^{*}}$. Additionally, $f$ has weak derivatives belonging to $L_{2}(\rho)$ for any probability measure $\rho$ on $\mathbb{R}^{d}$.
The proof can be found in Appendix A.1.1. This lemma indicates that the space of functions $\mathcal{F}_{\infty}$ is contained within the space of $1/2$-Hölder continuous functions with weak derivatives that are square integrable with respect to any probability measure. Recall that on a compact, all Lipschitz functions are Hölder continuous functions, indicating that the Hölder condition is less restrictive.
Now, we consider the relationship of $\mathcal{F}_{\infty}$ to other function spaces. Starting with the one-dimensional case, BKerNN reduces to kernel ridge regression with the Brownian kernel, which is also equivalent to learning with natural cubic splines (for an introduction to splines, see Wahba, 1990). For multi-dimensional data, we use the Fourier decomposition of functions to bound the defining norms of function spaces, enabling us to make comparisons.
@@
Assume we only consider functions $f$ with support on the ball centred at 00 with radius $R$ and norm $\|\cdot\|^{*}$. Assume $f$ has a Fourier decomposition,
@@
The proof is given in Appendix A.1.2. We remark that the condition $\int_{\mathbb{R}^{d}}|\hat{f}(\omega)|\cdot\|\omega\|\,{\rm d}\omega<\infty$ is a form of constraint on the regularity of the first-order derivatives.
@@ 2.4 Learning the Kernel or Training a Neural Network?
We first transform the optimisation problem before considering our setup from two different perspectives: one through kernel learning and the other through neural networks. To transform the optimisation problem, we use the representer theorem, a well-known result in RKHS that allows us to replace the optimisation over functions in the RKHS with optimisation over a finite weighted sum of the kernel at the data points.
@@
with $K=\int_{\mathbb{R}^{d}}K^{(w)}\,{\rm d}\nu(w)$ and and $K^{(w)}\in\mathbb{R}^{n\times n}$ is the kernel matrix for kernel $k^{(B)}$ and data $(w^{\top}x_{1},\ldots,w^{\top}x_{n})$. At the optimum, the support of $\mu$ from Equation (4) can be obtained from that of $\nu$ from Equation (9) by normalising all vectors.
Notice that there is shift in spaces, as $\nu$ is a probability distribution on $\mathbb{R}^{d}$, whereas $\mu$ was a probability distribution on $\mathcal{S}^{d-1}$. The proof is provided in Appendix A.2.2.

### v2

@@ Abstract
We propose a new method for feature learning and function estimation in supervised learning via regularised empirical risk minimisation. Our approach considers functions as expectations of Sobolev functions over all possible one-dimensional projections of the data. This framework is similar to kernel ridge regression, where the kernel is $\mathbb{E}_{w}(k^{(B)}(w^{\top}x,w^{\top}x^{\prime}))$, with $k^{(B)}(a,b):=\min(|a|,|b|)\mathds{1}_{ab>0}$ the Brownian kernel, and the distribution of the projections $w$ is learnt. This can also be viewed as an infinite-width one-hidden layer neural network, optimising the first layer’s weights through gradient descent and explicitly adjusting the non-linearity and weights of the second layer. We introduce a gradient-based computational method for the estimator, called Brownian Kernel Neural Network (BKerNN), using particles to approximate the expectation, where the positive homogeneity of the Brownian kernel leads to improved robustness to local minima. Using Rademacher complexity, we show that BKerNN’s expected risk converges to the minimal risk with explicit high-probability rates of $O(\min((d/n)^{1/2},n^{-1/6}))$ (up to logarithmic factors). Numerical experiments confirm our optimisation intuitions, and BKerNN outperforms kernel ridge regression, and favourably compares to a one-hidden layer neural network with ReLU activations in various settings and real data sets.
@@ 1 Introduction
Multiple index models have been extensively studied, leading to various methods for estimating the feature space. Brillinger (2012) introduced the method of moments for Gaussian data and one feature, by using specific moments to eliminate the unknown function. For features of any dimension, several methods have been proposed. Sliced inverse regression (SIR) (Li, 1991) uses second-order moments to identify effective dimensions by slicing the response variable and finding linear combinations of predictors, while improvements have been proposed (Yang et al., 2017), these methods heavily rely on assumptions about the covariate distribution shape and prior knowledge of the distribution. Iterative improvements have been an interesting line of work (Dalalyan et al., 2008), while optimisation-based methods like local averaging minimise an objective function to estimate the subspace (Fukumizu et al., 2009; Xia et al., 2002). Despite their practical performance, particularly the MAVE method (Xia et al., 2002), the theoretical guarantees show exponential dependence on the original data dimension, making them less suitable for high-dimensional settings. More recently, neural networks have also been applied to this problem, as demonstrated by Mousavi-Hosseini et al. (2024) and Bietti et al. (2023) (for the single-index model), who both focused on the continuous limit of the optimisation process. Furthermore, Bach (2017) has shown that one-hidden-layer neural networks of infinite width with ReLU activation are adaptive to hidden linear features. Unlike earlier methods, these neural network approaches simultaneously learn both the feature space and the prediction function, a key difference from the other presented methods. Extensions have also been proposed to tackle non-linear feature learning as in the introductory work by Moniri et al. (2024).
@@
Regularising the empirical risk minimization objective with a penalty term helps guide estimator behaviour (towards few relevant features for example) and can often be interpreted as a form of function space regularisation. Recent advances in norm-based capacity measures in neural networks have deepened the understanding of how such regularisation influences generalisation. In particular, the path norm, originally introduced for ReLU networks, provides strong generalization guarantees due to its scale-invariant properties (Liu et al., 2024). This notion has been refined through the basis-path norm, which disentangles independent basis paths, offering a more precise characterization of neural network complexity (Zheng et al., 2018). Building on these ideas, Parhi and Nowak (2020) extend the analysis from a variational perspective, linking path norm regularisation to Banach space representations of neural networks, while Galanti et al. (2023) focus on norm-based bounds for sparse networks.
Now, in the specific context of feature learning (and not necessarily neural networks), Rosasco et al. (2013) used derivatives for regularisation in nonparametric models, focusing on variable selection. While their method reduces to classical regularisation techniques for linear functions, it faces limitations: functions depending on a single variable do not belong in the chosen RKHS, using derivatives at data points limits the exploitation of regularity, and there is no benefit from hidden variables in the misspecified case. An improvement over this method was studied for both feature learning and variable selection by Follain and Bach (2024), where a trace norm penalty (Koltchinskii et al., 2011; Giraud, 2014) on the derivatives was used for the feature learning case. However, the dependency on the dimension of the rate did not allow high-dimensional learning. We can justify the use of trace norm penalties by considering the structure of neural networks. Under the multiple-index model, the weights $w_{1},\ldots,w_{m}$ of the first layer are expected to lie in a low-rank subspace of rank at most $k$. However, directly enforcing a rank constraint is not practical for optimisation. Therefore, we could use a relaxation such as $\Omega(f)=\operatorname*{tr}\big{(}\big{(}\sum_{j=1}^{m}w_{j}w_{j}^{\top}\big{)}^{1/2}\big{)}$, which is the trace norm of a matrix containing the weights, to approximate the rank constraint effectively. However, there is still the issue of multiple constant terms (often referred to as “biases”) for a single weight. We will see specialised penalties for feature learning for a different family of functions.
@@ 2.1 Custom Space of Functions
Definition 1 (Infinite-Width Function Space)
@@
:=$\mathcal{F}_{\infty}:=\left\{f\mid f(\cdot)=c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu(g,w),\quad\Omega_{0}(f)<\infty\right\},$
@@
$\Omega_{0}(f):=\inf_{\mu\in\mathcal{P}(\mathcal{H}\times\mathcal{S}^{d-1})}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}\,{\rm d}\mu(g,w),$ (2)
with the infimum over $\mu$ such that $f=c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot)\,{\rm d}\mu(g,w)$, where $\|g\|_{\mathcal{H}}^{2}:=\int_{-\infty}^{+\infty}(g^{\prime})^{2}$.
Note that $c$ corresponds to the value of the function $f$ evaluated at the null vector and is therefore unique, which is not necessarily the case of the measure $\mu$. $\mathcal{F}_{\infty}$ is well-defined using a “variation norm” on couples $(g,w)$ integrated w.r.t a Borel measure on ×$\mathcal{H}\times\mathcal{S}^{d-1}$. This is the variation norm associated to the map $(g,w)\in\mathcal{H}\times\mathcal{S}^{d-1}\to g(w^{\top}\cdot)$ (which is a function from $\mathbb{R}^{d}$ to $\mathbb{R}$, see Kurkova and Sanguineti (2001) and Bach (2024, Section 9.3.2).
The function space $\mathcal{F}_{\infty}$ is inspired by infinite-width single hidden layer neural networks: with the addition of the intercept $c$, each function in this space can be seen as the integral of a linear part $w$ and a non-linearity ${\color[rgb]{0,0,0}g}$ over some probability distribution, as in Equation (1) where the non-linearity is $\eta\sigma(\cdot)$. Thus here the activation functions are learnt.
@@
We now consider regularised empirical risk minimisation starting with the basic penalty $\Omega_{0}$. This penalty enforces the regularity of the function and, because we use penalisation with non-squared norms, limits the cardinality of the support of $\mu$ on the space $\mathcal{H}$. While this penalty is not specifically aimed at feature learning, by limiting the number of non-zero particles, it indirectly promotes feature learning to some extent. This serves as a starting point, and we introduce more targeted penalties in Section 2.5 with a stronger feature learning behaviour. For $f\in\mathcal{F}_{m}$ written as in Definition 2, the penalty simplifies to $\Omega_{0}(f)=\frac{1}{m}\sum_{j=1}^{m}\|g_{j}\|_{\mathcal{H}}$. The learning objective is thus defined as
@@
$\min_{c\in\mathbb{R},{\color[rgb]{0,0,0}\mu\in\mathcal{P}(\mathcal{H}\times\mathcal{S}^{d-1})}}\frac{1}{n}\sum_{i=1}^{n}\ell\bigg{(}y_{i},c+\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}g(w^{\top}x_{i}){\rm d}{\color[rgb]{0,0,0}\mu(g,w)}\bigg{)}+\lambda\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\|g\|_{\mathcal{H}}{\rm d}{\color[rgb]{0,0,0}\mu(g,w)},$ (4)
@@
$\min_{c\in\mathbb{R},w_{1},\ldots,w_{m}\in\mathcal{S}^{d-1},g_{1},\ldots,g_{m}\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\ell\bigg{(}y_{i},c+\frac{1}{m}\sum_{j=1}^{m}g_{j}(w_{j}^{\top}x_{i})\bigg{)}+\lambda\frac{1}{m}\sum_{j=1}^{m}\|g_{j}\|_{\mathcal{H}},$ (5)
@@ 2.2 Properties of Reproducing Kernel Hilbert Space $\mathcal{H}$ and Kernel $k$
with the inner product $\langle\tilde{g},g\rangle_{\mathcal{H}}=\int_{\mathbb{R}}\tilde{g}^{\prime}g^{\prime}$. This space is a reproducing kernel Hilbert space with the reproducing kernel $k^{(B)}(a,b)=(|a|+|b|-|a-b|)/2=\min(|a|,|b|)\mathds{1}_{ab>0}$. This kernel, which can be referred to as the “Brownian” kernel, corresponds to the covariance of the Brownian motion at times $a$ and $b$ (Mishura and Shevchenko, 2017). To see that $k^{(B)}$ is the reproducing kernel of $\mathcal{H}$, it suffices to check that we have the reproducing property, i.e. that
$\forall g\in\mathcal{H},\forall a\in\mathbb{R},g(a)=\langle g,k^{(B)}_{a}\rangle_{\mathcal{H}},$
where $k^{(B)}_{a}:b\in\mathbb{R}\to k^{(B)}(a,b)\in\mathbb{R}$ [^2] Indeed, we have for $g\in\mathcal{H}$, and without loss of generality $a\in\mathbb{R}^{+}$ that $\langle g,k^{(B)}_{a}\rangle_{\mathcal{H}}=\int_{\mathbb{R}}g^{\prime}(b)(k^{(B)}_{a})^{\prime}(b)\ {\rm d}b=\int_{0}^{a}g^{\prime}(b)\ {\rm d}b=g(a)-g(0)=g(a)$ As a reproducing kernel, it is positive definite, meaning that for any $n\in\mathbb{N}$, $\alpha\in\mathbb{R}^{n}$, and $a\in\mathbb{R}^{n}$, we have $\sum_{i,j=1}^{n}\alpha_{i}k^{(B)}(a_{i},a_{j})\alpha_{j}\geq 0$. Additionally, we observe that $\|k^{(B)}_{a}\|^{2}_{\mathcal{H}}=\langle k^{(B)}_{a},k^{(B)}_{a}\rangle=k^{(B)}(a,a)=|a|$ and we can obtain $\|k^{(B)}_{a}-k^{(B)}_{b}\|^{2}_{\mathcal{H}}=|a-b|$ in a similar manner. It is also noteworthy that by definition, the functions in $\mathcal{H}$ are necessarily continuous, in fact even $1/2$-Hölder continuous as we see in Lemma 3.
@@ 2.3 Characterisation of $\mathcal{F}_{\infty}$
$\mathcal{F}_{\infty}$ is a vector space and $\max(f(0),\Omega_{0}(f))$ is a norm on $\mathcal{F}_{\infty}$. For $f\in\mathcal{F}_{\infty}$, the function $f$ is $1/2$-Hölder continuous with constant $\Omega_{0}(f)$, i.e., $|f(x)-f(x^{\prime})|\leq\Omega_{0}(f)\sqrt{\|x-x^{\prime}\|^{*}}$.
The proof can be found in Appendix A.1.1. This lemma indicates that the space of functions $\mathcal{F}_{\infty}$ is contained within the space of $1/2$-Hölder continuous functions. Recall that on a compact, all Lipschitz functions are Hölder continuous functions, indicating that the Hölder condition is less restrictive.
Now, we consider the relationship of $\mathcal{F}_{\infty}$ to other function spaces. Starting with the one-dimensional case, BKerNN reduces to kernel ridge regression with the Brownian kernel, which is also equivalent to learning with natural cubic splines (for an introduction to splines, see Wahba, 1990). Thus, in dimension one, we can compare precisely our function space (square-integrable derivative), to the function space associated with neural networks (Heeringa et al., 2024, Theorem 1), with rectified linear units (integrable second-derivative), and with sign activation function (integrable first-derivative).
For multi-dimensional data, we use the Fourier decomposition of functions to bound the defining norms of function spaces, enabling us to make comparisons.
@@
We consider a function $f$ with support on the ball centred at 00 with radius $R$ and norm $\|\cdot\|^{*}$ (Note that we do not assume that $f\in\mathcal{F}_{\infty}$). Assume $f$ has a Fourier transform and can be written using the inverse transform [^3] A sufficient condition is that both $f$ and $\hat{f}$ belong to $L^{1}(\mathbb{R}^{d})$. as
@@
The proof is given in Appendix A.1.2. We remark that the condition $\int_{\mathbb{R}^{d}}|\hat{f}(\omega)|\cdot\|\omega\|\,{\rm d}\omega<\infty$ is a form of constraint on the regularity of the first-order derivatives (as the Fourier transform of the gradient of $f$ is equal to $i\omega$ times the Fourier transform of $f$).
@@ 2.4 Learning the Kernel or Training a Neural Network?
We first transform the optimisation problem before considering our setup from two different perspectives: one through kernel learning and the other through neural networks. To transform the optimisation problem, we use the representer theorem, a well-known result in RKHS theory that allows us to replace the optimisation over functions in the RKHS with optimisation over a finite weighted sum of the kernel at the data points.
@@
with $K=\int_{\mathbb{R}^{d}}K^{(w)}\,{\rm d}\nu(w)$ and and $K^{(w)}\in\mathbb{R}^{n\times n}$ is the kernel matrix for kernel $k^{(B)}$ and data $(w^{\top}x_{1},\ldots,w^{\top}x_{n})$.
Notice that there is a shift in spaces, as $\nu$ is a probability distribution on $\mathbb{R}^{d}$, whereas $\mu$ was a probability distribution on ×${\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}$. The proof is provided in Appendix A.2.2.
In summary, we aim to perform multiple kernel learning with a kernel parametrized by a probability measure (Lemma 6), and approximate it using particles (Lemma 5).